OpenStudy (anonymous):

Use the point-slope form of the linear equation to find an equation of the line with the given slope and passing through the given point. Then write the equation in standard form. Slope 4; through (1, 9)

OpenStudy (saifoo.khan):

y = mx + b 9 = 4 + c c = 5 ===================================== y = mx + b y = 4x + 5

OpenStudy (angela210793):

y-y1=k(x-x1) whr (x1;y1) r the coordiantes of the point given and k= slope y-9=4(x-1) y-9=4x-4 y=4x+5

OpenStudy (anonymous):

a. 4x - y = -9 b. 4x - y = 35 c. 4x - y = -5 d. 4x - y = -13

OpenStudy (angela210793):

c. 4x - y = -5